OpenStudy (anonymous):

What number is 45% of 680?

OpenStudy (anonymous):

use ratios. 45/100 = x/680. Cross multiply and solve

OpenStudy (anonymous):

0.45 times 680 =306

OpenStudy (anonymous):

.45*689=answer

OpenStudy (anonymous):

@caroleen yup that's what I got too.